alter working 4 hours, frank has made 21 machine parts. at the same rate, how many parts can he make in 7 hours?
Found 2 solutions by troyapplehelen, checkley71:
Answer by troyapplehelen(46)

1. 21machines/4 hrs---21:4
2. x machines/7 hrs---x:7
3. the same rate---21:4=x:7, 4x=147, x=36.7 so, about 37 machines.

Answer by checkley71(8403)

4/7=21/X OR 4X=147 OR X=147/4 OR X=36.75